### v1

@@ 2 Preliminaries and main resuls
where $U_{i}=(u_{i},v_{i},\theta_{i},q_{i})\in\mathcal{H},\;i=1,2$. Define an operator $\mathcal{A}:{\cal D}({\cal A})\subseteq\mathcal{H}\to\mathcal{H}$ as
@@
$\mathcal{D}(\mathcal{A})=\left\{(u,v,\theta,q)\in\mathcal{H}\,\left|\,v\in\mathcal{D}(A^{1\over 2}),\;\sigma Au-A^{\alpha}\theta\in\mathcal{D}(A^{-\frac{\gamma}{2}}),\;-A^{\alpha}v+A^{\frac{\beta}{2}}q\in H,\;q+A^{\frac{\beta}{2}}\theta\in H\right.\right\}.$
@@
Assume that there exists $U_{n}=(u_{n},v_{n},\theta_{n},q_{n})^{\top}\in\mathcal{D}(\cal A)$ such that $U_{n}\to U_{0}:=(u_{0},v_{0},\theta_{0},q_{0})^{\top}$ in $\mathcal{H}$, $\mathcal{A}U_{n}\to W:=(w_{1},w_{2},w_{3},w_{4})^{\top}$ in $\mathcal{H}$, i.e.,
@@
It suffices to show $U_{0}\in\mathcal{D}(\cal A)$ and $\mathcal{A}U_{0}=W$.
@@
Since $U_{n}\in\mathcal{D}(\cal A)$, which implies $-q_{n}-A^{\beta/2}\theta_{n}\in H$, then $B_{n}:=-A^{-\beta/2}q_{n}-\theta_{n}\in\mathcal{D}(A^{\beta/2})$. Recalling that $A$ is a self-adjoint, positive definite operator on $H$, $q_{n}\to q_{0}$ in $H$, therefore, $A^{-\beta/2}q_{n}\to A^{-\beta/2}q_{0}$ in $H$, which together with $\theta_{n}\to\theta_{0}$ in $H$ implies
@@
We have shown in Lemma 2.1 that $\mathcal{A}$ is a closed operator. Note that $\mathcal{D}(\mathcal{A})\supseteq{\cal D}_{0}:=\mathcal{D}(A^{1-\frac{\gamma}{2}})\times\mathcal{D}(A^{\alpha})\times\mathcal{D}(A^{\alpha-\frac{\gamma}{2}})\times\mathcal{D}(A^{\frac{\beta}{2}})$, and ${{\overline{{\cal D}_{0}}}=\mathcal{H}}$, thus $\mathcal{A}$ is densely defined on $\mathcal{H}$. Furthermore, for any vector $U=(u,v,\theta,q)^{\top}\in\mathcal{D}(\mathcal{A})$, we have
@@
indicating that $\mathcal{A}$ is dissipative. By direct calculation, one gets $\mathcal{A}^{*}:\mathcal{D}(\mathcal{A}^{*})\to\mathcal{H}$ is dissipative as well. By [15, Corollary I.4.4], $\mathcal{A}$ is the generator of a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ of contraction on $\mathcal{H}$. ∎
@@
In this paper, we explore the long-time behavior of the solution $T(\cdot)U_{0}$ of system (2.3) for initial values $U_{0}$ belonging to $\mathcal{D}({\mathcal{A}})\cap\mathcal{R}(\mathcal{A})$. Note that since $\mathcal{A}$ generates a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ of contractions on $\mathcal{H}$, $-\mathcal{A}$ is a sectorial operator and thereby $\mathcal{R}({\mathcal{A}}^{s}(I-{\mathcal{A}})^{-(s+t)})=\mathcal{R}({\mathcal{A}}^{s})\cap\mathcal{D}({\mathcal{A}}^{t})$ for $s,t\geq 0$ by [3, Proposition 3.10]. The main results are stated as follows.
@@
Let $m>0$ and $(\alpha,\beta,\gamma)\in Q$, $\mathcal{H}$ and $\mathcal{A}$ be defined by (2.1) and (2.2), respectively. Suppose that $(T(t))_{t\geq 0}$ is the $C_{0}$-semigroup of contractions generated by ${\mathcal{A}}$. Then as $t\to\infty,$
@@
Let $m=0$ and $(\alpha,\beta)\in Q^{*}$, $\mathcal{H}$ and $\mathcal{A}$ be defined by (2.1) and (2.2), respectively. Suppose that $(T(t))_{t\geq 0}$ is the $C_{0}$-semigroup of contractions generated by $\mathcal{A}$. Then
$\big{\|}T(t){\mathcal{A}}(I-{\mathcal{A}})^{-2}\big{\|}=O(t^{-\frac{\alpha}{2\alpha-\beta}}),\quad t\to\infty.$ (2.18)
@@
At the end of this section, we present the following interpolation lemma, which will be utilized in subsequent sections.
@@
Let $A\,:\,{\cal D}(A)\subseteq H$ be self-adjoint and positive definite, $r,\;p\;,q\in\mathbb{R}$. Then
$\|A^{p}x\|\leq\|A^{q}x\|^{p-r\over q-r}\|A^{r}x\|^{q-p\over q-r},\quad\forall\;r\leq p\leq q,\;x\in{\cal D}(A^{q}).$ (2.21)
@@
$\begin{array}[]{ll}&\|s(isI-\mathcal{A})^{-1}\|=O(1),\;\;s\to 0,\\ &\|\lambda^{-k}(i\lambda I-\mathcal{A})^{-1}\|=O(1),\;\;\lambda\to\infty.\end{array}$ (3.3)
@@
In summary, by (3.14), (3.16), (3.18) and (3.22), we conclude $\|U_{1,n}\|_{\mathcal{H}}=\|(u_{1,n},v_{1,n},\theta_{1,n},q_{1,n})\|_{\mathcal{H}}=o(1)$.
@@
${{\cal R}e}(\sigma\theta_{2,n},\,A^{1-\alpha}u_{2,n})-{{\cal R}e}{(q_{2,n},\,A^{\frac{\beta}{2}-\alpha}(I+mA^{\gamma})v_{2,n})}-\|\theta_{2,n}\|^{2}+\|v_{2,n}\|^{2}+m\|A^{\frac{\gamma}{2}}v_{2,n}\|^{2}=\lambda_{n}^{-k}o(1).$ (3.27)
@@
$\lambda_{n}^{-1}A^{\alpha-\frac{\beta}{2}}v_{2,n}=-{\color[rgb]{1.00,0.00,0.00}\definecolor[named]{pgfstrokecolor}{rgb}{1.00,0.00,0.00}i}A^{-\frac{\beta}{2}}\theta_{2,n}+\lambda_{n}^{-1}q_{2,n}+\lambda_{n}^{-k-1}o(1){.}$ (3.32)
@@
Using the same method in [6], we know that the eigenvalues $\lambda_{n}$ of operator ${\cal A}$ satisfy the following quartic equation:
@@
where $\mu_{n},\;n=1,2,\cdots$ are the eigenvalues of the operator $A$. By [6, Section 5.1], we get that the solutions to (3.34) when $(\alpha,\beta,\gamma)\in Q$ are as follows:
@@
In summary, by (4.9), (4.10), (4.12) and (4.15), we obtain $\|U_{1,n}\|_{\mathcal{H}}=\|(u_{1,n},v_{1,n},\theta_{1,n},q_{1,n})\|_{\mathcal{H}}=o(1)$.
@@
We see the first term of (4.19) tends to 0 because of (4.17) and (4.18), the second term tends to 0 because of (4.16). Thus,
@@
Finally, we shall prove the decay order is sharp by a similar argument as in Section 3. Note that $k=\frac{2\alpha-\beta}{\alpha}\geq 1$ always holds. Suppose $\mu_{n},\;\lambda_{n},\;n=1,2,\cdots$ are the eigenvalues of operators $A$ and ${\cal A}$, respectively. Then we have the following quartic equation using the same method in [6, Section 5],

### v2

@@ 2 Preliminaries and main resuls
where $U_{i}=(u_{i},v_{i},\theta_{i},q_{i})^{\top}\in\mathcal{H},\;i=1,2$. Define an operator $\mathcal{A}:{\mathcal{D}}(\mathcal{A})\subseteq\mathcal{H}\to\mathcal{H}$ as
@@
$\mathcal{D}(\mathcal{A})=\left\{(u,v,\theta,q)^{\top}\in\mathcal{H}\,\left|\,\begin{array}[]{l}v\in\mathcal{D}(A^{1\over 2}),\;\sigma Au-A^{\alpha}\theta\in\mathcal{D}(A^{-\frac{\gamma}{2}}),\\ -A^{\alpha}v+A^{\frac{\beta}{2}}q\in H,\;q+A^{\frac{\beta}{2}}\theta\in H\end{array}\right.\right\}.$
@@
Assume that there exists $U_{n}=(u_{n},v_{n},\theta_{n},q_{n})^{\top}\in\mathcal{D}(\mathcal{A})$ such that $U_{n}\to U_{0}:=(u_{0},v_{0},\theta_{0},q_{0})^{\top}$ in $\mathcal{H}$, $\mathcal{A}U_{n}\to W:=(w_{1},w_{2},w_{3},w_{4})^{\top}$ in $\mathcal{H}$, i.e.,
@@
It suffices to show $U_{0}\in\mathcal{D}(\mathcal{A})$ and $\mathcal{A}U_{0}=W$.
@@
Since $U_{n}\in\mathcal{D}(\mathcal{A})$, which implies $-q_{n}-A^{\beta/2}\theta_{n}\in H$, then $B_{n}:=-A^{-\beta/2}q_{n}-\theta_{n}\in\mathcal{D}(A^{\beta/2})$. Recalling that $A$ is a self-adjoint, positive definite operator on $H$, and $q_{n}\to q_{0}$ in $H$; therefore, $A^{-\beta/2}q_{n}\to A^{-\beta/2}q_{0}$ in $H$, which together with $\theta_{n}\to\theta_{0}$ in $H$ implies
@@
We have shown in Lemma 2.1 that $\mathcal{A}$ is a closed operator. Note that $\mathcal{D}(\mathcal{A})\supseteq\mathcal{D}_{0}:=\mathcal{D}(A^{1-\frac{\gamma}{2}})\times\mathcal{D}(A^{\alpha})\times\mathcal{D}(A^{\alpha-\frac{\gamma}{2}})\times\mathcal{D}(A^{\frac{\beta}{2}})$, and ${{\overline{\mathcal{D}_{0}}}=\mathcal{H}}$, thus $\mathcal{A}$ is densely defined on $\mathcal{H}$. Furthermore, for any vector $U=(u,v,\theta,q)^{\top}\in\mathcal{D}(\mathcal{A})$, we have
@@
indicating that $\mathcal{A}$ is dissipative. By direct calculation, one gets $\mathcal{A}^{*}:\mathcal{D}(\mathcal{A}^{*})\to\mathcal{H}$ is dissipative as well. By [18, Corollary I.4.4], $\mathcal{A}$ is the generator of a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ of contraction on $\mathcal{H}$. ∎
@@
In this paper, we explore the long-time behavior of the solution $T(\cdot)U_{0}$ of system (2.3) for initial values $U_{0}$ belonging to $\mathcal{D}({\mathcal{A}})\cap\mathcal{R}(\mathcal{A})$. Note that since $\mathcal{A}$ generates a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ of contractions on $\mathcal{H}$, $-\mathcal{A}$ is a sectorial operator and thereby $\mathcal{R}({\mathcal{A}}^{s}(I-{\mathcal{A}})^{-(s+t)})=\mathcal{R}({\mathcal{A}}^{s})\cap\mathcal{D}({\mathcal{A}}^{t})$ for $s,t\geq 0$ by [3, Proposition 3.10]. The main results are stated as Theorems 2.5 and 2.7.
@@
Let $m>0$ and $(\alpha,\beta,\gamma)\in Q$, $\mathcal{H}$ and $\mathcal{A}$ be defined by (2.1) and (2.2), respectively. Suppose that $(T(t))_{t\geq 0}$ is the $C_{0}$-semigroup of contractions generated by ${\mathcal{A}}$. Then for $t\geq 1,$
@@
Let $m=0$ and $(\alpha,\beta)\in Q^{*}$, $\mathcal{H}$ and $\mathcal{A}$ be defined by (2.1) and (2.2), respectively. Suppose that $(T(t))_{t\geq 0}$ is the $C_{0}$-semigroup of contractions generated by $\mathcal{A}$. Then for $t\geq 1,$
$\big{\|}T(t){\mathcal{A}}(I-{\mathcal{A}})^{-2}\big{\|}=O(t^{-\frac{\alpha}{2\alpha-\beta}}).$ (2.18)
@@
At the end of this section, we present the following interpolation lemma (see, e.g., [7] and [15]), which will be utilized in subsequent sections.
@@
Let $A\,:\,\mathcal{D}(A)\subseteq H$ be self-adjoint and positive definite, $r,\;p\;,q\in\mathbb{R}$. Then
$\|A^{p}x\|\leq\|A^{q}x\|^{p-r\over q-r}\|A^{r}x\|^{q-p\over q-r},\quad\forall\;r\leq p\leq q,\;x\in\mathcal{D}(A^{q}).$ (2.21)
Remark 2.10.
The study of stability of abstract thermoelastic system (1.1) has been extensively covered in [7, 9, 10] and this paper. In [9], the exponential stability region of system (1.1) was obtained. However, later in [10], Han, Kuang, and Zhang demonstrated that the exponentially stable region could be extended if $\sigma\tau=1$. Furthermore, they investigated polynomial stability of system (1.1) when $(\alpha,\beta,\gamma)\in Q^{c}:=[0,1]\times[0,1]\times(0,1]\setminus Q$ taking certain specific values. Following this track, [7] provided a detailed analysis of the polynomial stability of system (1.1) for every $(\alpha,\beta,\gamma)\in Q^{c}$. Finally, this paper considers the polynomial stability of system (1.1) when $(\alpha,\beta,\gamma)\in Q$.
In all the aforementioned papers, the proof techniques are similar, utilizing frequency domain methods combined with interpolation Lemma 2.9 and detailed spectral analysis. The key difference is that for the case $(\alpha,\beta,\gamma)\in Q^{c}$, the results were proved by contradiction based on [4], which requires $i\mathbb{R}\subseteq\rho(\mathcal{A})$. However, when $(\alpha,\beta,\gamma)\in Q$, since 0 is in the spectrum of the generator $\mathcal{A}$, we use Lemma 2.8 instead of [4] to prove the result. Another difference is that the method used to prove the well-posedness of system (1.1) in this paper (Theorem 2.4) is relatively new.
@@
$\begin{array}[]{ll}\|s(isI-\mathcal{A})^{-1}\|&=O(1),\;\;s\to 0,\\ \|\lambda^{-k}(i\lambda I-\mathcal{A})^{-1}\|&=O(1),\;\;\lambda\to\infty.\end{array}$ (3.3)
@@
In summary, by (3.14), (3.16), (3.18) and (3.22), we conclude
$\|U_{1,n}\|_{\mathcal{H}}=\|(u_{1,n},v_{1,n},\theta_{1,n},q_{1,n})\|_{\mathcal{H}}=o(1).$
@@
$\begin{array}[]{lll}&{{\cal R}e}(\sigma\theta_{2,n},\,A^{1-\alpha}u_{2,n})-{{\cal R}e}{(q_{2,n},\,A^{\frac{\beta}{2}-\alpha}(I+mA^{\gamma})v_{2,n})}\\ &\quad-\|\theta_{2,n}\|^{2}+\|v_{2,n}\|^{2}+m\|A^{\frac{\gamma}{2}}v_{2,n}\|^{2}=\lambda_{n}^{-k}o(1).\end{array}$ (3.29)
@@
$\lambda_{n}^{-1}A^{\alpha-\frac{\beta}{2}}v_{2,n}=-iA^{-\frac{\beta}{2}}\theta_{2,n}+\lambda_{n}^{-1}q_{2,n}+\lambda_{n}^{-k-1}o(1){.}$ (3.34)
@@
The eigenvalues $\lambda$ of operator ${\cal A}$ satisfy the following quartic equation:
$\begin{array}[]{lll}f(\lambda,\mu_{n})&\equiv(m\tau\cdot\mu_{n}^{\gamma}+\tau)\lambda^{4}+(m\cdot\mu_{n}^{\gamma}+1)\lambda^{3}\\ &\quad+(\tau\cdot\mu_{n}^{2\alpha}+m\cdot\mu_{n}^{\beta+\gamma}+\sigma\tau\cdot\mu_{n}+\mu_{n}^{\beta})\lambda^{2}\\ &\quad+(\mu_{n}^{2\alpha}+\sigma\cdot\mu_{n})\lambda+\sigma\cdot\mu_{n}^{1+\beta}\\ &=0.\end{array}$ (3.36)
Without loss of generality, choose $\sigma=2$ and $\tau=1$. Taking $m=1$ we derive the characteristic equation of the system from (3.36):
@@
where $\lambda_{n}$ is the eigenvalue of $\mathcal{A}$ associated with $\mu_{n}\rightarrow\infty$ as $n\rightarrow\infty$. We aim to solve (3.37) in the setting $\mu_{n}\rightarrow\infty$. Using a similar method as shown in [7, 8, 10, 12], we get that the solutions to (3.37) when $(\alpha,\beta,\gamma)\in Q$ are as follows:
@@
In summary, by (4.9), (4.10), (4.12) and (4.15), we obtain
$\|U_{1,n}\|_{\mathcal{H}}=\|(u_{1,n},v_{1,n},\theta_{1,n},q_{1,n})\|_{\mathcal{H}}=o(1).$
@@
We see the first term of (4.19) tends to 0 because of (4.17) and (4.18), and the second term tends to 0 because of (4.16). Thus,
@@
Finally, we shall prove the decay order is sharp by a similar argument as Section 3. Note that $k=\frac{2\alpha-\beta}{\alpha}\geq 1$ always holds. Suppose $\mu_{n},\;\lambda_{n},\;n=1,2,\cdots$ are the eigenvalues of operators $A$ and ${\cal A}$, respectively. Then repeat the process at the end part of Section 3, we have the following quartic equation by taking $m=0,\sigma=2,\tau=1$ in (3.36),